# LeanKAN: A Parameter-Lean Kolmogorov-Arnold Network Layer with Improved Memory Efficiency and Convergence Behavior

### v1

@@ V Conclusions
In this work we proposed LeanKAN as an alternative to MultKAN for incorporating multiplication nodes into KAN layers. While MultKAN is not suitable for general use in a multidimensional output layer, LeanKAN’s reversed structure lends it to global use in output layers (at the cost of its lack of use for input layers). LeanKAN is also parameterized more compactly than MultKAN and is inherently less complex with fewer activations and fewer parameters for a given number of layers, nodes, and gridpoints. In a similar vein, it eliminates one of two multiplication hyperparameters, simplifying the network tuning process.
@@
LeanKAN is a one-to-one replacement for standard AddKAN or MultKAN layers, lending it to use not just in standard KAN applications, but also in augmented network structures based on KAN. In the two case studies presented here, we demonstrate LeanKAN’s performance both as a standard KAN layer in a toy multiplication reconstruction case, and also as the backbone of a KAN-ODE. Across the board, we found that LeanKAN’s compact structure did not carry any performance downgrades. In fact, LeanKAN was found to significantly outperform similarly sized and even substantially larger MultKANs in all studied metrics, thanks to its elimination of dummy activations and extraneous parameters. For general use in KANs with multiplication-based layers, we hope these results inspire adoption of LeanKANs as a simpler, leaner, and faster layer structure.
@@
The LeanKAN code will be made public upon acceptance of the manuscript.

### v2

@@ 5 Conclusions
In this work we proposed LeanKAN as an alternative to MultKAN for incorporating multiplication nodes into KAN layers. While MultKAN is not suitable for general use in a multidimensional output layer, LeanKAN’s reversed structure lends it to global use in output layers (at the cost of a restriction for input layers). LeanKAN is also parameterized more compactly than MultKAN and is inherently less complex with fewer activations and fewer parameters for a given number of layers, nodes, and gridpoints. In a similar vein, it eliminates one of two multiplication hyperparameters, simplifying the network tuning process.
@@
LeanKAN is a one-to-one replacement for standard AddKAN or MultKAN layers, lending it to use not just in standard KAN applications, but also in augmented network structures based on KAN. We demonstrate these capabilities in three case studies of increasing complexity, from a toy multiplication case to a complex-valued PDE. Across the board, we found that LeanKAN’s compact structure did not lead to any performance downgrades. In fact, LeanKAN was found to significantly outperform similarly sized and even substantially larger MultKANs and AddKANs in all studied metrics, thanks to its elimination of dummy activations and extraneous parameters. For general use in KANs with multiplication-based layers, we hope these results inspire adoption of LeanKANs as a simpler, leaner, and faster layer structure.
@@
The LeanKAN code is available at the following link: https://github.com/DENG-MIT/LeanKAN.
@@ Acknowledgement
Appendix A. Noisy and Irregular Data: Lotka-Volterra
[…]